# Automata Learning from Preference and Equivalence Queries

### v2

@@ Appendix A Technical Appendix
Lemma 5.
Suppose $\langle S,E,T;\mathcal{C},\Gamma\rangle$ is a unified, closed, and consistent symbolic observation table. If the hypothesis $\hat{h}=\langle\hat{Q},\Sigma^{I},\Sigma^{O},\hat{q}_{0},\hat{\delta},\hat{L}\rangle$ from MakeHypothesis$(\langle S,E,T;\mathcal{C},\Gamma\rangle)$ has $n$ states, and $\hat{h}$ was generated using a satisfying solution $\Lambda$ to constraints $\mathcal{C}$, then any other Moore machine $h=\langle Q,\Sigma^{I},\Sigma^{O},q_{0},\delta,L\rangle$ with $n$ or fewer states that is also consistent with $\Lambda$ and $T$ is isomorphic to $\hat{h}$.
Proof.
We make additions and appropriate adjustments to the proof of Angluin (1987). First, let us partition $\hat{Q}$ into $k=|\Sigma^{O}|$ disjoint sets $\hat{F}_{1},...,\hat{F}_{k}$ such that for all $\sigma_{j}\in\Sigma^{O}$, $\hat{q}\in\hat{F}_{j}$ if and only if $\hat{L}(\hat{q})=\sigma_{j}$. Note that it is possible for some of the $\hat{F}_{j}$ to be empty. Similarly, let us also partition $Q$ into $k$ disjoint sets $F_{1},...,F_{k}$ such that for all $\sigma_{j}\in\Sigma^{O}$, $q\in F_{j}$ if and only if $L(q)=\sigma_{j}$. Next, we can define for all $q\in Q$ and for all $\sigma_{j}\in\Sigma^{O}$ the function $\textbf{row}^{\wedge}(q):E\rightarrow\Sigma^{O}$ such that $\textbf{row}^{\wedge}(q)(e)=\sigma_{j}$ if and only if $\delta(q,e)\in F_{j}$.
@@
Next, we can consider bijective mappings between $\hat{Q}$ and $Q$. Since $Q$ and $\hat{Q}$ have the same cardinality, and because $\textbf{row}^{\wedge}(\delta(q_{0},s))\equiv\textbf{row}(s)$, we know that for every $s\in S$ corresponds to a unique $q\in Q$, specifically, $q=\delta(q_{0},s)$. We can define the bijective mapping $\textbf{row}^{-\wedge}:\hat{Q}\rightarrow Q$ via for all $s\in S$, $\textbf{row}^{-\wedge}(\textbf{row}(s))=\delta(q_{0},s)$. From this mapping, we observe that $\textbf{row}^{-\wedge}(\textbf{row}(\varepsilon))=\textbf{row}^{-\wedge}(\hat{q}_{0})=\delta(q_{0},\varepsilon)=q_{0}$. Additionally, for all $s\in S$ and for all $\sigma\in\Sigma$, $\textbf{row}^{-\wedge}(\hat{\delta}(\textbf{row}(s),\sigma))=\textbf{row}^{-\wedge}(\textbf{row}(s\cdot\sigma))=\delta(q_{0},s\cdot\sigma)=\delta(\delta(q_{0},s),\sigma)=\delta(\textbf{row}^{-\wedge}(\textbf{row}(s)),\sigma)$, which implies that for all $s\in S$ and $\sigma\in\Sigma$, $\textbf{row}^{-\wedge}(\hat{\delta}(\textbf{row}(s),\sigma))=\delta(\textbf{row}^{-\wedge}(\textbf{row}(s)),\sigma)$.
@@
Lemma 6.
Suppose $\langle S,E,T;\mathcal{C},\Gamma\rangle$ is a unified, closed, and consistent symbolic observation table. Let $n$ be the number of states in the hypothesis MakeHypothe-sis$(\langle S,E,T;\mathcal{C},\Gamma\rangle)$. Any Moore machine consistent with $\Gamma$ and $\mathcal{C}$ must have at least $n$ states.
Proof.
@@
Proof.
Lemma 1 establishes that all hypotheses in $\mathcal{H}$ have equivalent states and equivalent transitions, which implies that every hypothesis in $\mathcal{H}$ has the same number of states. Lemmas 2 and 4 together establish that in a hypothesis, if a specific sequence $s$ is classified correctly, then all other sequences which start at the same initial state as $s$ and end in the same state as $s$ will be classified the same as $s$. Thus, if a state is classified correctly, then all sequences terminating at that state are classified correctly. Finally, Lemma 5 establishes that any Moore machine $d$ consistent with $\Lambda$ and $T$ from $\hat{h}=\textsc{MakeHypothesis}(\langle S,E,T;\mathcal{C},\Gamma\rangle)$ must be isomorphic to $\hat{h}$; otherwise $d$ must contain at least one more state than $\hat{h}$. This implies that $\hat{h}$ is the smallest Moore machine consistent with $\Lambda$ and $T$, and because $\Lambda$ is taken from all possible satisfying solutions to $\mathcal{C}$, then by Lemma 1, every hypothesis $h\in\mathcal{H}$ consistent with $\mathcal{C}$ has that same smallest number of states. Therefore, any other Moore machine not in $\mathcal{H}$, but still consistent with $\mathcal{C}$, must have more states. ∎
@@
Theorem 3 (Query Complexity).
@@
Definition 6 (Probably Approximately Correct Identification).
@@
Theorem 4.
Remap achieves probably approximately correct identification of any Moore machine when the teacher $\mathcal{T}$ uses sampling-based equivalence queries.

### v3

@@ 0.A.6 Technical Proofs
Proposition 5
Suppose $\langle S,E,T;\mathcal{C},\Gamma\rangle$ is a unified, closed, and consistent symbolic observation table. If the hypothesis $\hat{h}=\langle\hat{Q},\Sigma^{I},\Sigma^{O},\hat{q}_{0},\hat{\delta},\hat{L}\rangle$ from the function MakeHypothesis$(\langle S,E,T;\mathcal{C},\Gamma\rangle)$ has $n$ states, and $\hat{h}$ was generated using a satisfying solution $\Lambda$ to constraints $\mathcal{C}$, then any other Moore machine $h=\langle Q,\Sigma^{I},\Sigma^{O},q_{0},\delta,L\rangle$ with $n$ or fewer states that is also consistent with $\Lambda$ and $T$ is isomorphic to $\hat{h}$.
Proof
We make additions and appropriate adjustments to the proof of Lemma 4 from Angluin [5]. First, let us partition $\hat{Q}$ into $k=|\Sigma^{O}|$ disjoint sets $\hat{F}_{1},...,\hat{F}_{k}$ such that for all $\sigma_{j}\in\Sigma^{O}$, $\hat{q}\in\hat{F}_{j}$ if and only if $\hat{L}(\hat{q})=\sigma_{j}$. Note that it is possible for some of the $\hat{F}_{j}$ to be empty. Similarly, let us also partition $Q$ into $k$ disjoint sets $F_{1},...,F_{k}$ such that for all $\sigma_{j}\in\Sigma^{O}$, $q\in F_{j}$ if and only if $L(q)=\sigma_{j}$. Next, we can define for all $q\in Q$ and for all $\sigma_{j}\in\Sigma^{O}$ the function $\textbf{row}^{\wedge}(q):E\rightarrow\Sigma^{O}$ such that $\textbf{row}^{\wedge}(q)(e)=\sigma_{j}$ if and only if $\delta(q,e)\in F_{j}$.
@@
Next, we can consider bijective mappings between $\hat{Q}$ and $Q$. Since $Q$ and $\hat{Q}$ have the same cardinality, and because $\textbf{row}^{\wedge}(\delta(q_{0},s))\equiv\textbf{row}(s)$, we know that for every $s\in S$ corresponds to a unique $q\in Q$, specifically, $q=\delta(q_{0},s)$. We can define the bijective mapping $\textbf{row}^{-\wedge}:\hat{Q}\rightarrow Q$ via for all $s\in S$, $\textbf{row}^{-\wedge}(\textbf{row}(s))=\delta(q_{0},s)$. From this mapping, we observe that $\textbf{row}^{-\wedge}(\textbf{row}(\varepsilon))=\textbf{row}^{-\wedge}(\hat{q}_{0})=\delta(q_{0},\varepsilon)=q_{0}$. Additionally, for all $s\in S$ and for all $\sigma\in\Sigma$, $\textbf{row}^{-\wedge}(\hat{\delta}(\textbf{row}(s),\sigma))=\textbf{row}^{-\wedge}(\textbf{row}(s\cdot\sigma))=\delta(q_{0},s\cdot\sigma)=\delta(\delta(q_{0},s),\sigma)=\delta(\textbf{row}^{-\wedge}(\textbf{row}(s)),\sigma)$, which implies for all $s\in S$ and $\sigma\in\Sigma$, $\textbf{row}^{-\wedge}(\hat{\delta}(\textbf{row}(s),\sigma))=\delta(\textbf{row}^{-\wedge}(\textbf{row}(s)),\sigma)$.
@@
Proposition 6
Suppose $\langle S,E,T;\mathcal{C},\Gamma\rangle$ is a unified, closed, and consistent symbolic observation table. Let $n$ be the number of states in the hypothesis returned from MakeHypothesis$(\langle S,E,T;\mathcal{C},\Gamma\rangle)$. Any Moore machine consistent with $\Gamma$ and $\mathcal{C}$ must have at least $n$ states.
Proof
@@
Proof
Proposition 1 establishes that all hypotheses in $\mathcal{H}$ have equivalent states and equivalent transitions, which implies that every hypothesis in $\mathcal{H}$ has the same number of states. Propositions 2 and 4 together establish that in a hypothesis, if a specific sequence $s$ is classified correctly, then all other sequences which start at the same initial state as $s$ and end in the same state as $s$ will be classified the same as $s$. Thus, if a state is classified correctly, then all sequences terminating at that state are classified correctly. Finally, Proposition 5 establishes that any Moore machine $d$ consistent with $\Lambda$ and $T$ from $\hat{h}=\textsc{MakeHypothesis}(\langle S,E,T;\mathcal{C},\Gamma\rangle)$ must be isomorphic to $\hat{h}$; otherwise $d$ must contain at least one more state than $\hat{h}$. This implies that $\hat{h}$ is the smallest Moore machine consistent with $\Lambda$ and $T$, and because $\Lambda$ is taken from all possible satisfying solutions to $\mathcal{C}$, then by Proposition 1, every hypothesis $h\in\mathcal{H}$ consistent with $\mathcal{C}$ has that same smallest number of states. Therefore, any other Moore machine not in $\mathcal{H}$, but still consistent with $\mathcal{C}$, must have more states.∎
@@
Theorem 0.A.3 (Query Complexity)
@@
Definition 12 (Probably Approximately Correct Identification)
@@
Theorem 0.A.4